OpenStudy (anonymous):

Why is 1^infinity not 1?

OpenStudy (anonymous):

It is 1...

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1 to the power of any number is 1...

OpenStudy (anonymous):

\[1^\infty\] is undefined

OpenStudy (paxpolaris):

huhh.... ???

OpenStudy (anonymous):

@ganeshie8

OpenStudy (paxpolaris):

if you are being nitpicky ..... Infinity is not a number .... you can't involve infinity in Math operations. you have to use limit to express what you are trying to say: \[\lim_{n \to \infty}1^n=1\]

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

@chiel140 see infinity is a concept and not a number. Is says that you give me a number, i can give you a bigger number, and it increases without any bounds. If it was a valid number, then the answer would be 1, but it is not a number and hence it is 1^ infinity undefined/. If someone asks you, what is 1^cow, can you say the answer? No, because it is not a number.

OpenStudy (anonymous):

ah, ok

OpenStudy (anonymous):

thanks :)

OpenStudy (perl):

1^oo can equal to different numbers, that is why it is indeterminate

OpenStudy (perl):

the limit is usually defined in a case by case basis, you have to do some kind of logarithm argument . but there are different limits

OpenStudy (perl):

@PaxPolaris i think the lim 1^n , as n-> oo is 1

OpenStudy (anonymous):

but, \[\lim_{x \rightarrow \infty}(1+1/x)^x=e\]

OpenStudy (perl):

right, thats why its indeterminate

OpenStudy (perl):

lim (1 + 1/x ) ^(2x) = e^2

OpenStudy (anonymous):

It's so hard to understand infinity

OpenStudy (perl):

1 / infinity = 0 , always , doesnt matter the type of infinity

OpenStudy (perl):

you just have to memorize a few indeterminate forms

OpenStudy (perl):

oo/oo , 0/0 , oo - oo , 0^0, and 1^oo

OpenStudy (anonymous):

what about oo^0 ?

OpenStudy (perl):

oo^0 doesn't make sense

OpenStudy (perl):

wait

OpenStudy (perl):

lim n ^(1/n) , n -> oo

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}x^{1/x}\]

OpenStudy (paxpolaris):

the limit of 1/n IS 1, correct.

OpenStudy (perl):

lets use n

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@chiel140 in this kind of situations we generally take logarithms and check if 'Hospital is applicable or not

OpenStudy (perl):

oo/oo , 0/0 , oo - oo , 0^0, 1^oo , oo^0

OpenStudy (perl):

now try to find me another one (i dare you )

OpenStudy (anonymous):

:p I don't think I can

OpenStudy (perl):

the rest of the combinations are not indeterminate

OpenStudy (perl):

oo/0 is not indeterminate, do you see why?

OpenStudy (paxpolaris):

a/0 ... ha!

OpenStudy (perl):

because oo/0 = oo * 1/0 = oo * (+/- oo )

OpenStudy (perl):

right 1/0 -> +/- oo ,

OpenStudy (perl):

oo/0 = oo * 1/0 = oo * (+/- oo ) = +/- oo^2 = +/- oo

OpenStudy (perl):

infinity * infinity = infinity infinity + infinity = infinity infinity - infinity = undefined

OpenStudy (anonymous):

yeah

OpenStudy (perl):

- infinity - infinity = -2* infinity = - infinity

OpenStudy (perl):

-infinity + infinity = undefined

OpenStudy (anonymous):

this is an infinite conversation.:P

OpenStudy (perl):

these are infinity rules for limits (it may not be valid in other contexts)

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} x ^{1/x}=e ^{\lim_{x \rightarrow \infty }\ln(x)/x}=e ^{\lim_{x \rightarrow \infty }1/x}=e^0=1\]

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

@Princer_Jones

OpenStudy (anonymous):

yes it is correct.

OpenStudy (anonymous):

:) thanks